OpenStudy (anonymous):

Find for solutions of -2x-y=-2. Write the solutions as ordered pairs.

OpenStudy (bibby):

solve into y = mx+b form \( -2x-y=-2\\ -2x=y-2\\ y=-2x+2\) now plug in random x's and solve for y's

OpenStudy (anonymous):

it says that I need 4 ordered pairs

OpenStudy (bibby):

plug in x=0,1,2,3 and solve